How do I plot this equation on a graph? y=|x|+3
Answer by Edwin McCravy(20086)

How do I plot this equation on a graph? y=|x|+3

Graphs of absolute value functions are V-shaped graphs.
We first find the vertex (the sharp point of the "V"}

1.  Set what's inside the absolute value bars = 0 and solve
    for x:

    x = 0

2.  Substitute that into the equation:
    
    y=|x|+3
    y=|0|+3
    y=0+3
    y=3

3.  Plot the vertex (0,3)



4.  Get a point on the right of 
    the vertex:

    Let x=1
    y=|x|+3
    y=|1|+3
    y=1+3
    y=4

    That's one point (1,4).  Plot it:



5.  Get a point on the left of 
    the vertex:

    Let x=-1
    y=|x|+3
    y=|-1|+3
    y=1+3
    y=4

    That's another point (-1,4).  Plot it:
   


Draw the V-shaped graph:


Edwin